OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable y is eliminated in the first and second equations, then the second and third equations. 3x – y + 3z = 1 4x + y – 6z = 3 4x – 2y + 3z = –3 7x – 3z = 4 12x – 9z = 3 7x – 3z = 4 4x – 6z = 3 3x + 3z = 1 12x – 9z = 3 3x + 3z = 1 4x – 6z = 3

Parth (parthkohli):

I prefer a question at a time. I'd help you with the first.

Parth (parthkohli):

Add all equations.

OpenStudy (anonymous):

Lol it is a question, 9ut the last 8 sets are choices.

OpenStudy (anonymous):

4 sets*

Parth (parthkohli):

lol

OpenStudy (anonymous):

Ha ha ha..

Parth (parthkohli):

Start by adding the top two equations.

OpenStudy (anonymous):

What do you get?? @Loujoelou

OpenStudy (anonymous):

I know it's either C/D, cause 3x+3z=1 is what you get when you eliminate y from 1st/2nd equation.

OpenStudy (anonymous):

nvm other way around. It's A/B

OpenStudy (anonymous):

Cause you get 7x-3z=4 not what I said earlier.

Parth (parthkohli):

Yep.

Parth (parthkohli):

Now, you multiply the second equation by 2 and then add equation 1 and 2.

OpenStudy (anonymous):

Now multiply second equation by 2 and then add second and third equation and let us see what you get.

OpenStudy (anonymous):

K I get it now :) It's A cause you get 12x-9z=3. Was just a little confused about it.

Parth (parthkohli):

Yep! :)

OpenStudy (anonymous):

tyvm @waterineyes @ParthKohli I understand it perfectly now :) At least doing this by elimination method I mean.

Parth (parthkohli):

You got the thing successfully. Congratulations, Lou!

OpenStudy (anonymous):

You are doing this by elimination only @Loujoelou

OpenStudy (anonymous):

Lol thx @ParthKohli and yes for this problem I had to @waterineyes

OpenStudy (anonymous):

Okay..